OpenStudy (anonymous):

Absolute Value Equations! Please need help totally lost! Solve for x: -3|x + 7| = -12 Answers: x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

Directrix (directrix):

-3 * |x + 7| = -12 Divde both sides by -3 and post what you get, okay?

Directrix (directrix):

@jblock21

OpenStudy (anonymous):

7/3 = 2.3? - 12 / 3 = - 4

Directrix (directrix):

You cannot get into the absolute value yet. Suppose you had this equation: -3 * y = -12 Divide by sides by -3 and post what you get.

OpenStudy (anonymous):

-3 would be 0 then -12 would be -9?

Directrix (directrix):

No.